OpenStudy (varmstrong6698):

∆ABC has side lengths of 10 units, 20 units, and 24 units. ∆XYZ is similar to ∆ABC, and the length of its longest side is 60 units. The perimeter of ∆XYZ is ? units. If the height of ∆ABC, with respect to its longest side being the base, is 8 units, the area of ∆XYZ is ? square units. first ? 108 110 135 270 Second ? 240 480 600 1200

OpenStudy (varmstrong6698):

Will fan and give medal!

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

We are told the longest side of the triangle XYZ measures 60. Since triangles ABC and XYZ are similar, then the longest side of triangle ABC must correspond to the longest side of triangle XYZ.

OpenStudy (varmstrong6698):

I think the first one is 135. is that correct?

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

\(\dfrac{60}{24} = \dfrac{P}{10 + 20 + 24} \) \(P = 135\)

OpenStudy (varmstrong6698):

Okay and is the second one 600?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

You are correct again.

OpenStudy (varmstrong6698):

Thank You (:

OpenStudy (mathstudent55):

You're welcome.